OpenStudy (anonymous):

what is the sine of pi/2

OpenStudy (klimenkov):

It equals 1.

OpenStudy (anonymous):

thanks

Parth (parthkohli):

In degree language,\[\sin\left({\pi \over 2}\right) = \sin(90^{\circ}) = 1\]

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

careful....

Parth (parthkohli):

 We observe that we are left with hypotenuse over hypotenuse because the opposite side to 90 is just the hypotenuse.